Can tyou help please with the equation: calculate the products: (4+3i)(7-3i)
Found 2 solutions by macston, lwsshak3:
Answer by macston(5194)
Answer by lwsshak3(11628)

Can tyou help please with the equation: calculate the products: (4+3i)(7-3i)
***
i=√-1
i^2=-1
i^3=-i
i^4=1
..
FOIL
(4+3i)(7-3i)
=28+21i-12i-9i^2
=28+9i+9
=37+9i